# IOVS4NeRF: Incremental Optimal View Selection for Large-Scale NeRFs

### v1

@@ V-D Ablations Study
  • Full Model: IOVS4NeRF with completed hybrid-uncertainty computation strategy shown in Eq. (6).
@@ V-E Benefit to Classical Photogrammetry Solutions
Note that the proposed approach IOVS4NeRF is a flexible framework to assist NeRF-based novel view synthesis. It is interesting if IOVS4NeRF can assist the classical photogrammetry pipeline. In this experiment, we use the selected views by IOVS4NeRF, random, FVS and completed image set Full as input to famous commercial photogrammetry software MestShape to perform the dense 3D reconstruction pipeline for comparison. Specifically, we randomly 25% images in Lake and Polytech datasets as initialization and select 25% views in the rest set by using IOVS4NeRF, random, FVS for incrementally re-training. The results in Figure. 9 show that the reconstructed scenes by using the selected views from random, FVS are with obvious artifacts. In contrast, the selected views by IOVS4NeRF lead to similar 3D reconstruction quality as the ones using Full image sets as input for MetaShape but with only half processing time. This experiment demonstrates IOVS4NeRF not only can ameliorate the novel view synthesis task by using NeRFs, but also benefit the classical photogrammetry solutions.
@@ VI Conclusion
In this work, we present a novel framework method, IOVS4NeRF, with the goal of rendering large-scale urban scenes. Our approach builds upon the original NeRF architecture by using Instant-NGP for assistance, which incorporates enhanced explicit-implicit fusion for information storage and lightweight neural MLP networks can accelerate the overall reconstruction speed. The threshold sampling with joint rendering of uncertainty and Voronoi graph perception can improve overall rendering quality, making them suitable for large-scale 3D reconstruction tasks.
When applied to extensive urban landscapes, our approach overcomes the limitations of existing methods. Our model achieves high visual fidelity rendering even in extremely large-scale city scenes, a critical capability for real-world applications. Empirical evaluations of our fast renderer based on our framework suggest that interactive large-scale rendering remains an open research challenge. Our model inherits some limitations from NeRF-based approaches, such as handling a large number of high-resolution images, indicating areas for further refinement in future work.

### v2

@@ V-D Ablation Study
  • Full Model: IOVS4NeRF with completed hybrid uncertainty computation strategy shown in Eq. (7).
@@ V-E Benefit to Classical Photogrammetry Solutions
Note that the proposed approach IOVS4NeRF is a flexible framework to assist NeRF-based novel view synthesis. It is interesting if IOVS4NeRF can assist the classical photogrammetry pipeline. In this experiment, we use the selected views by IOVS4NeRF, Random, FVS and completed image set Full as input to famous commercial photogrammetry software MetaShape to perform the dense 3D reconstruction pipeline for comparison. Specifically, we randomly 25% images in Lake and Polytech datasets as initialization and select 25% views in the rest set by using IOVS4NeRF, Random, FVS for incrementally re-training. The results in Figure. 9 show that the reconstructed scenes by using the selected views from Random, FVS are with obvious artifacts. In contrast, the selected views by IOVS4NeRF lead to similar 3D reconstruction quality as the ones using Full image sets as input for MetaShape but with only half processing time. This experiment demonstrates IOVS4NeRF not only can ameliorate the novel view synthesis task by using NeRFs, but also benefit the classical photogrammetry solutions.
@@ VI Conclusion
In conclusion, this study introduces IOVS4NeRF, an incremental optimal view selection framework aimed at addressing the limitations of NeRF in resource-constrained scenarios. By actively expanding the training dataset with strategically selected samples based on a novel hybrid uncertainty metric, IOVS4NeRF significantly enhances the quality of novel view synthesis while minimizing additional resource requirements. The efficacy of this approach has been validated through extensive experiments on realistic scenes, demonstrating its robustness even with limited training data. Future work will focus on further refining the model and exploring its applicability in various real-world scenarios. The implementation code will be released following the paper’s acceptance, facilitating further research and development in this domain.